OpenStudy (anonymous):

3logx(6)+4logx(6)=10

OpenStudy (jamesj):

First 3log_x(6) = log_x(6^3) and 4log_x(6) = log_x(6^4). So 3logx(6)+4logx(6) = log_x(6^3) + log_x(6^4) = log_x(6^3 . 6^4) = log_x(6^7) Hence if 3logx(6)+4logx(6)=10 then log_x(6^7) = 10 Or in other words, 6^7 = x^10. I.e., x = 6^(7/10) = ....

OpenStudy (jamesj):

@Hero: Actually no -- you've transcribed the equation into Wolfram incorrectly. The x here is not the argument of the log function, but its base.